How do you find surface area if the rectangle is 100 feet by 200 feet with quarter circles in the corners that have a radius of 20 feet.
Answer by richwmiller(17219)

You did NOT say what you are trying to find the surface area of.
the quarter circles, the rectangle, the rectangle minus the quarter circles

Find the area of the rectangle. 100*200=20,000 ft^2
Call it A(R).
Since there are 4 quarters it is one entire circle
area of the circle is pi*r^2
A(C)=3.14*20^2=3.14*400=1256
if you want the rectangle minus the quarter circles
now A(R)-A(C)=20,000 -1256=